40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 281 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  2745   Wed Mar 31 19:29:58 2010 HartmutUpdateElectronics(1cm-) Si PD transfer functions update

Recorded transfer functions for the 1cm Si-PD as described on p. 2708

for different biases. I put the plots in there, to keep the info in one place,

where the label on the PD case (which Steve made without asking him) points

to.

I talked to some people recently about the fact that the responsivity (A/W) of the PD

changes even at DC for different biases. I tested this again and should be more precise about this:

The first time I observed this was in the transfer functions as shown on p. 2708.

With 'DC' I meant 'low frequency' there, as you can still see an effect of the bias as low as 100kHz.

Then at one point I saw the responsivity changing with bias also at true DC.

However, it turned out that this is only the case if the photocurrent is too high.

If the photocurrent is 4mA, you need 400mV bias to get the max. responsivity.

For 2mA photocurrent, the responsivity is already maximal for 0V bias.

An effect for relative low frequencies remains however.

The DC check of responsivity was done with white light from a bulb.

 

 

  2744   Wed Mar 31 16:55:05 2010 josephbUpdateComputers2 computers from Alex and Rolf brought to 40m

I went over to Downs today and was able to secure two 8 core machines, along with mounting rails.  These are very thin looking 1U chassis computers. I was told by Rolf the big black box computers might be done tomorrow afternoon.  Alex also kept one of the 8 core machines since he needed to replace a hard drive on it, and also wanted to keep for some further testing, although he didn't specify how long.

I also put in a request with Alex and Rolf for the RCG system to produce code which includes memory location hooks for plant models automatically, along with a switch to flip from the real to simulated inputs/outputs.

 

  2743   Wed Mar 31 16:31:44 2010 steveUpdatePEMGuralp interface box turned off

Quote:

Quote:

I went and double-checked and aligned the styrofoam cooler at ~5:00 UTC. It was fine, but we really need a better huddling box. Where's that granite anyway?

Here's the new Huddle Test output. This time I show the X-axis since there's some coherence now below 0.1 Hz.

You'll also notice that the Wiener filter is now beating the FD subtraction. This happened when I increased the # of taps to 8000. Looks like the noise keeps getting lower as I increase the number of taps, but this is really a kind of cheat if you think about it carefully.

 The same thing happening again.  The intermittent offset upstream of the seismometer that never got fixed.

The granite plate and ball bearings are in. I will place seismometers on it.

 

Attachment 1: grlpntrfcbxoff.jpg
grlpntrfcbxoff.jpg
  2742   Wed Mar 31 15:31:53 2010 steveUpdatePSLReference Cavity RF PD base upgraded

Quote:

Some more words about the RFAM: I noticed that there was an excess RFAM by unlocking the RC and just looking at the RF out with the 50 Ohm input of the scope. It was ~100 mVp-p! In the end our method to minimize the AM was not so sensible - we aligned the waveplate before the EOM so as to minimize the p-pol light transmitted by the PBS cube just ahead of the AOM. At first, this did not minimize the RFAM. But after I got angry at the bad plastic mounting of the EOM and re-aligned it, the AM seemed to be small with the polarization aligned to the cube. It was too small to measure on the scope and on the spectrum analyzer, the peak was hopping around by ~10-20 dB on a few second timescale. Further reduction would require some kind of active temperature stabilization of the EOM housing (maybe a good SURF project!).

For the EOM mount we (meaning Steve) should replace the lame 2-post system that's in there with one of the mounts of the type that is used in the Mach-Zucker EOMs. I think we have spare in the cabinet next to one of the arms.

After the RFAM monkeying, I aligned the beam to the RC using the standard, 2-mirror, beam-walking approach. You can see from the attached plot that the transmission went up by ~20% ! And the reflection went down by ~30%. I doubt that I have developed any new alignment technique beyond what Yoichi and I already did last time. Most likely there was some beam shape corruption in the EOM, or the RFAM was causing us to lock far off the fringe. Now the reflected beam from the reference cavity is a nice donut shape and we could even make it better by doing some mode matching! This finally solves the eternal mystery of the bad REFL beam (or at least sweeps it under the rug).

At the end, I also fixed the alignment of the RFPD. It should be set so the incident angle of the beam is ~20-40 deg, but it was instead set to be near normal incidence ?! Its also on flimsy plastic legs. Steve, can you please replace this with the new brass ones?

 Teflon feet removed and heavy brass-delrin pd base installed. Ref-cavity reflected light remains to be beautiful doughnut shape on camera.

Attachment 1: brspdbs.JPG
brspdbs.JPG
  2741   Wed Mar 31 12:30:31 2010 ranaUpdateGreen Lockingfrequency counter for green PLL

Its a good measurement - you should adjust the input range of the 620 using the front panel 'scale' buttons to see how the noise compares to Matt's circuit when the range is reduced to 1 MHz. In any case, we would use it in the 350 MHz range mode. What about the noise of the frequency discriminator from MITEQ?

  2740   Wed Mar 31 11:52:32 2010 kiwamuSummaryGreen LockingRe:conversion efficiency of PPKTP

Good point. There is a trick  to avoid a divergence.

Actually the radius of the cylindrical wave is set to the spot size at the surface of the crystal instead of an actual beam waist. This is the idea Dmass told me before.

So that the radius is expressed by w=w0(1+(L/2ZR)2)1/2, where w0 is beam waist, L is the length of the crystal and ZR is the rayleigh range.

In this case the radius can't go smaller than w0/2 and the solution can not diverge to infinity.

Quote:

Question:

Why does the small spot size for the case (A) have small efficiency as the others? I thought the efficiency goes diverged to infinity as the radius of the cylinder gets smaller.

 

 

 

  2739   Wed Mar 31 10:34:02 2010 josephbUpdateelogElog not responding this morning

When I went to use the elog this morning, it wasn't responding.  I killed the process on nodus, and then restarted, per the 40m wiki instructions.

  2738   Wed Mar 31 03:45:49 2010 MottHowToComputersNew script for controlling the AG4395A

 

I took data for the 2 NPRO PLL using the script I wrote for the AG4395, but it is very noisy above about 1 MHz.  I assume this is something to do with the script (since I am fairly confident we don't have 600 dB response in the PZT), so I will go in tomorrow to more carefully understand what is going on, I may need to include a bit more latency in the script to allow the NA to settle a bit more.  I am attaching the spectrum just to show the incredibly high noise level, 

Attachment 1: noisy_spec.png
noisy_spec.png
  2737   Wed Mar 31 02:57:48 2010 kiwamuUpdateGreen Lockingfrequency counter for green PLL

Rana found that we had a frequency counter SR620 which might be helpful for lock acquisition of the green phase lock.

It has a response of 100MHz/V up to 350MHz which is wide range and good for our purpose. And it has a noise level of 200Hz/rtHz @ 10Hz which is 1000 times worse than that Matt made (see the entry).

The attached figure is the noise curve measured while I injected a signal of several 100kHz. In fact I made sure that the noise level doesn't depends on the frequency of an input signal.

The black curve represents the noise of the circuit Matt has made, the red curve represents that of SR620.

Attachment 1: FCnoise.png
FCnoise.png
  2736   Tue Mar 30 22:13:49 2010 KojiSummaryGreen Lockingconversion efficiency of PPKTP

Question:

Why does the small spot size for the case (A) have small efficiency as the others? I thought the efficiency goes diverged to infinity as the radius of the cylinder gets smaller.

Quote:

With a 30mm PPKTP crystal the conversion efficiency from 1064nm to 532nm is expected to 3.7 %/W.

Therefore we will have a green beam of more than 2mW by putting 700mW NPRO.

Last a couple of weeks I performed a numerical simulation for calculating the conversion efficiency of PPKTP crystal which we will have.

Here I try to mention about just the result. The detail will be followed later as another entry.


The attached figure is a result of the calculation.

The horizontal axis is the waist of an input Gaussian beam, and the vertical axis is the conversion efficiency.

You can see three curves in the figure, this is because I want to double check my calculation by comparing  analytical solutions.

The curve named (A) is one of the simplest solution, which assumes that the incident beam is a cylindrical plane wave.

The other curve (B) is also analytic solution, but it assumes different condition; the power profile of incident beam is a Gaussian beam but propagates as a plane wave.

The last curve (C) is the result of my numerical simulation. In this calculation a focused Gaussian beam is injected into the crystal.

The numerical result seems to be reasonable because the shape and the number doesn't much differ from those analytical solutions.

 

  2735   Tue Mar 30 21:11:42 2010 kiwamuSummaryGreen Lockingconversion efficiency of PPKTP

With a 30mm PPKTP crystal the conversion efficiency from 1064nm to 532nm is expected to 3.7 %/W.

Therefore we will have a green beam of more than 20mW by putting 700mW NPRO.

Last a couple of weeks I performed a numerical simulation for calculating the conversion efficiency of PPKTP crystal which we will have.

Here I try to mention about just the result. The detail will be followed later as another entry.


The attached figure is a result of the calculation.

The horizontal axis is the waist of an input Gaussian beam, and the vertical axis is the conversion efficiency.

You can see three curves in the figure, this is because I want to double check my calculation by comparing  analytical solutions.

The curve named (A) is one of the simplest solution, which assumes that the incident beam is a cylindrical plane wave.

The other curve (B) is also analytic solution, but it assumes different condition; the power profile of incident beam is a Gaussian beam but propagates as a plane wave.

The last curve (C) is the result of my numerical simulation. In this calculation a focused Gaussian beam is injected into the crystal.

The numerical result seems to be reasonable because the shape and the number doesn't much differ from those analytical solutions.

Attachment 1: efficiency_waist_edit.png
efficiency_waist_edit.png
  2734   Tue Mar 30 11:16:05 2010 josephbHowToComputersezca update information (CDS SVN)

I'd like to try installing an updated multi-threaded ezca extension later this week, allowing for 64-bit builds of GDS ezca tools, provided by Keith Thorne.  The code can be found in the LDAS CVS under gds, as well as in CDS subversion repository, located at 

https://redoubt.ligo-wa.caltech.edu/websvn/

Its under gds/epics/ in that repository.  The directions are fairly simple:

1) to install ezca with mult-threading in an existing EPICS installation
-copy ezca_2010mt.tar.gz (EPICS_DIR)/extensions/src
-cd (EPICS_DIR)/extensions/src
-tar -C -xzf ezca_2010mt
-modify (EPICS_DIR)/extensions/Makefile to point 'ezca' at 'ezca_2010mt'
-cd ezca_2010mt
-set EPICS_HOST_ARCH appropriately
-make


 

 

  2733   Tue Mar 30 06:37:32 2010 ranaConfigurationPSLReference Cavity PD Noise Spectrum

Some more words about the RFAM: I noticed that there was an excess RFAM by unlocking the RC and just looking at the RF out with the 50 Ohm input of the scope. It was ~100 mVp-p! In the end our method to minimize the AM was not so sensible - we aligned the waveplate before the EOM so as to minimize the p-pol light transmitted by the PBS cube just ahead of the AOM. At first, this did not minimize the RFAM. But after I got angry at the bad plastic mounting of the EOM and re-aligned it, the AM seemed to be small with the polarization aligned to the cube. It was too small to measure on the scope and on the spectrum analyzer, the peak was hopping around by ~10-20 dB on a few second timescale. Further reduction would require some kind of active temperature stabilization of the EOM housing (maybe a good SURF project!).

For the EOM mount we (meaning Steve) should replace the lame 2-post system that's in there with one of the mounts of the type that is used in the Mach-Zucker EOMs. I think we have spare in the cabinet next to one of the arms.

After the RFAM monkeying, I aligned the beam to the RC using the standard, 2-mirror, beam-walking approach. You can see from the attached plot that the transmission went up by ~20% ! And the reflection went down by ~30%. I doubt that I have developed any new alignment technique beyond what Yoichi and I already did last time. Most likely there was some beam shape corruption in the EOM, or the RFAM was causing us to lock far off the fringe. Now the reflected beam from the reference cavity is a nice donut shape and we could even make it better by doing some mode matching! This finally solves the eternal mystery of the bad REFL beam (or at least sweeps it under the rug).

At the end, I also fixed the alignment of the RFPD. It should be set so the incident angle of the beam is ~20-40 deg, but it was instead set to be near normal incidence ?! Its also on flimsy plastic legs. Steve, can you please replace this with the new brass ones?

Attachment 1: rc.png
rc.png
  2732   Mon Mar 29 21:43:27 2010 AlbertoConfigurationPSLReference Cavity PD Noise Spectrum

[Rana, Alberto]

This evening we measured the noise spectrum of the reference cavity PD used in the FSS loop. From that we estimated the transimpedance and found that the PD is shot-noise limited. We also found a big AM oscillation in correspondence of the FSS modulation sideband which we later attenuated at least in part.

This plot shows the spectrum noise from the RF output of the photodetector.
 
 (here you should be able to see an attached figure, if not it's probably becasue imagemagic has having problems with displaying png files)
2010-03-29_FSS_PD_shotnoise_and_darknoise.png
 
The tall peak at 21.5 MHz is the AM modulation introduced by the EOM. It seems to be caused by a misalignment of the EOM which might be somehow modulating the polarization.
The mount in which the EOM sits is not very solid. We should change it with something similar to that of the other two EOMs in the Mach Zehnder.
By tightening down the plastic screws of the mount Rana reduced the amplitude of the AM modulation by 20dB.
 
The bump in both the dark and shot noise are in corrispondence of the resonance of the PD's electronics. As it appears, the electronics is not well tuned: the bump should coincide with the AM peak.
 
In the case of the dark noise spectrum, the bump is due to the thermal noise of the electronics. It's a good sign: it means that the electronics is good enough to be sensitive to it.
 
Transimpedance Estimate
As a "sanity check" we made an approximate estimate of the transimpedance to make sure that the PD is dominated by shot noise rather than other noises, ie electronic's noise.
 
  1. Supposing that the laser beam hitting the PD was shot noise limited, we measured 1.1V at the DC output. That let us estimate the photocurrent at DC of 20mA, for a 50Ohm output impedance.
  2. The shot noise for 20mA is 80 pA/rtHz
  3. From the nosie spectrum, we measured 3e-7 v/rtHz at 21.5MHz
  4. The impedance at RF is then Z_rf = 3e-7 V/rtHz / 80e-12 pA ~ 4000 Ohm
  5. Since the RF path inside the PD has a gain of 10, the transimpedance is ~400Ohm, which is about as we (ie Rana) remembered it to be.
  6. The PD seems to be working fine.
Attachment 2: 2010-03-29_FSS_PD_shotnoise_and_darknoise.png
2010-03-29_FSS_PD_shotnoise_and_darknoise.png
  2731   Mon Mar 29 18:50:14 2010 KojiUpdateSUSPRM sidemagnet glued

PRM was released from the fixuture without any trouble. This was the last magnet gluing until ETMs are delivered.

The below is the up-to-date Jenne stat table.

The clean room is getting too narrow. I am thinking that we should install ITMs to the chamber so that we can accommodate SRM/PRM suspensions.

Attachment 1: StatusTable.png
StatusTable.png
  2730   Mon Mar 29 18:41:34 2010 KojiConfigurationSUSStarted to build TTs

Steve and Koji

WE started to build 5 TTs. 4 of them are used in the recycling cavities. One is the spare.

We built the structure and are building the cantilever springs.

Attachment 1: IMG_2348.jpg
IMG_2348.jpg
  2729   Mon Mar 29 15:26:47 2010 MottHowToComputersNew script for controlling the AG4395A

I just put a script in the /cvs/cds/caltech/scripts/general/netgpibdata/ directory to control the network analyzer called AG4395A_Run.py .   A section has been added to the wiki with the other GPIB script sections (http://lhocds.ligo-wa.caltech.edu:8000/40m/netgpib_package#AG4395A_Run.py)

  2728   Mon Mar 29 15:19:33 2010 mevansUpdateGreen Lockingfrequency discriminator for green PLL

Thanks for the great entry!

In order to make this work for higher frequencies, I would add Hartmut's suggestion of a frequency dividing input stage.  If we divide the input down by 100, the overall range will be about 200MHz, and the noise will be about 20Hz/rtHz.  That might be good enough... but we can hope that the commercial device is lower noise!

Quote:

Last Friday, Matt made a frequency discriminator circuit on a bread board in order to test the idea and study the noise level. I think it will work for phase lock acquisition of Green locking.

As a result a response of 100kHz/V and a noise level of 2uV/rtHz @ 10Hz are yielded. This corresponds to 0.2Hz/rtHz @ 10Hz.

The motivation of using frequency discriminators is that  it makes a frequency range wider and easier for lock acquisition of PLLs in green locking experiment.

FD.png

  2727   Mon Mar 29 10:40:59 2010 josephbUpdateCamerasGigE camera no work from screen

Quote:

Not that this is an urgent concern, just a data point which shows that it doesn't just not work at the sites.

I had to restart the dhcpd server on Ottavia that allows us to talk to the camera.  I then also changed the configuration script on the camera so that it no longer thinks ottavia is 131.215.113.97, but correctly 192.168.113.97.  Overall took 5 minutes.

I also looked up services for Centos 5, and set it using the program serviceconf to start the DHCP server  when Ottavia is rebooted now.  That should head off future problems of that nature.  For reference, to start the dhcp server manually, become root and type "service dhcpd start".

 

  2726   Mon Mar 29 02:07:50 2010 KojiSummaryPSLFSS Work from Sunday: Open Loop Gain

Quote:

I measured the open loop gain of the FSS (as usual, I have multiplied the whole OLG by 10dB to account for the forward loop gain in the box). I used a source level of -20 dBm and made sure this was not saturating by changing the level.

Its clear that the BW is limited by the resonance at ~1.7 MHz. Does anyone know what that is?

 EO resonance in the RC path?

  2725   Mon Mar 29 01:45:26 2010 daisukeConfigurationGeneralPeriscope version B for green laser ...

Here the design of the periscope for the 90 deg steering version is posted.

straight version http://nodus.ligo.caltech.edu:8080/40m/2709

Attachment 1: 40m_periscope_B.png
40m_periscope_B.png
Attachment 2: 40m_periscope_B_100329.pdf
40m_periscope_B_100329.pdf 40m_periscope_B_100329.pdf 40m_periscope_B_100329.pdf 40m_periscope_B_100329.pdf
Attachment 3: 40m_periscope_B_dwg_100329.zip
  2724   Mon Mar 29 01:11:33 2010 ranaSummaryPSLFSS Work from Sunday: RF Out Spectrum

I measured the RF spectrum coming out the FSS RFPD to look for saturations - its close to the hairy edge. This is with the 8x power increase from my AOM drive increase. I will increase the FSS's modulation frequency which will lower the Q and gain of the PD to compensate somewhat. The lower Q will also gain us phase margin in the FSS loooop.

 

I put in a bi-directional 20 dB coupler (its only rated down to 30 MHz, but its only off by ~0.3 dB at 21 MHz) between the RFPD and the FSS box. I looked at the time series on the 300 MHz scope and measured the power spectrum.

The peak signal on the scope was 40 mV; that translates to 400 mV at the RFPD output. Depending on whether the series resistor in the box is 20 or 50 Ohms, it means the MAX4107 is close to saturating.

As you can see from the spectrum, its mostly likely to hit its slew rate limit (500 V/us) first. Actually its not going to hit the limit: but even getting within a factor of 10 is bad news in terms of distortion.

Besides the multiples of the modulation frequency, you can see that most of the RMS comes from the strange large peaks at 137.9 and 181.1 MHz. Anyone know what these are from?

TEK00000.PNGTEK00001.PNGTEK00002.PNG

On the middle plot above, I have enabled the 20 MHz BW limit so you can see how much the amplitude goes down when only the 21.5 MHz SB is included. You can also see from the leftmost plot that once in awhile there is some 400mV/10ns slewing. Its within a factor of 10 of the slew rate limit.

Attachment 1: rfout.png
rfout.png
  2723   Sun Mar 28 23:47:47 2010 ranaSummaryPSLFSS Work from Sunday: Open Loop Gain

I measured the open loop gain of the FSS (as usual, I have multiplied the whole OLG by 10dB to account for the forward loop gain in the box). I used a source level of -20 dBm and made sure this was not saturating by changing the level.

Its clear that the BW is limited by the resonance at ~1.7 MHz. Does anyone know what that is?

Attachment 1: fssloop.png
fssloop.png
Attachment 2: sweep2.png
sweep2.png
  2722   Sun Mar 28 23:17:46 2010 ranaSummaryPSLFSS Work from Sunday: noise spec

This is the error point spectrum - it is filled with huge multiples of ~75 kHz as Yoichi noticed a couple years ago.

I tried to use the netgpib.py package to read out the Agilent 4395, but the SVN had been corrupted by someone saving over the netgpib.py package. To get it to work on rosalba I reverted to the previous version, but whoever is busy hacking on netgpib.py needs to checkin the original package and work on some test code instead.

I also noticed that the default output format for the AG4395.py file is in units of Watts. This is kind of dumb - we need someone to develop this package a little as Yoichi did for the SRS785.

Attachment 1: in2.png
in2.png
  2721   Sun Mar 28 20:51:31 2010 ranaSummaryPSLFSS Work from Sunday: Cavity Suspension is Ridiculously Undamped!

What is the Transfer Function of the suspension of the reference cavity? What were the design requirements? What is the Q and how well does the eddy current damping work? What did Wolfowitz know about the WMD and when? Who cooked the RTV in there and why didn't we use Viton??

To get to the bottom of these questions, today I shook the cavity and measured the response. To read out the pitch and yaw modes separately, I aligned the input beam to be misaligned to the cavity. If the beam is mis-aligned in yaw, for example, the transmitted light power becomes first order sensitive to the yaw motion of the cavity.

In the attached image (10 minute second-trend), you can see the second trends for the transmitted and relfected power. The first ringdown comes from the pitch or vertical mode. The second (shorter) one comes from the yaw misalignment and the yaw shake.

To achieve the up/down shake, I leaned onto the table and pumped it at its eigenfrequency. For the yaw shake, I put two fingers on the RC can's sweater and pushed with several pounds of force at the yaw eigenfrequency (2.6 Hz). For the vertical, I jumped up and down at half the vertical eigenfrequency (4 Hz).

I also made sure that the .SCAN field on these EPICS records were set to 9 so that there is no serious effect from a beating between the eigenfrequency and the EPICS sample rate.

Punchline:

f_vert   = 4 Hz

tau_vert = 90 seconds

Q_vert   = 1000            (yes, that number over there has 3 zeros)

 

f_hor    = 2.6 Hz

tau_hor  = 30 seconds

Q_hor    = 250

 

This is an absurd and probably makes us very sensitive to seismic noise - let's make sure to open up the can and put some real rubber in there to damp it. My guess is that these high Q modes

are just the modes of the last-stage steel spring / pendulum.

Attachment 1: Untitled.png
Untitled.png
  2720   Sun Mar 28 20:05:33 2010 ranaSummaryPSLFSS Work from Sunday: AOM/VCO level set wrong

Just before working on the FSS today, I noticed that the VCO RF output level was set incorrectly.

This should ALWAYS be set so as to give the maximum power in the first order diffracted sideband. One should set this by maximizing the out of lock FSS RFPD DC level to max.

The value was at 2.8 on the VCOMODLEVEL slider. In the attached plot (taken with the FSS input disabled) you can see that this puts us in the regime where the output power to the FSS is first order sensitive to the amplitude noise on the electrical signal to the AOM. This is an untenable situation.

For adjusting the power level to the FSS, we must always use the lamba/2 plate between the AOM and the RC steering mirrors. This dumps power out to the side via a PBS just before the periscope.

Attachment 1: Untitled.png
Untitled.png
  2719   Sun Mar 28 20:00:17 2010 ranaUpdateCamerasGigE camera no work from screen

Not that this is an urgent concern, just a data point which shows that it doesn't just not work at the sites.

Attachment 1: Untitled.png
Untitled.png
  2718   Sun Mar 28 17:28:26 2010 matt, kiwamuUpdateGreen Lockingfrequency discriminator for green PLL

Last Friday, Matt made a frequency discriminator circuit on a bread board in order to test the idea and study the noise level. I think it will work for phase lock acquisition of Green locking.

As a result a response of 100kHz/V and a noise level of 2uV/rtHz @ 10Hz are yielded. This corresponds to 0.2Hz/rtHz @ 10Hz.

The motivation of using frequency discriminators is that  it makes a frequency range wider and easier for lock acquisition of PLLs in green locking experiment.

For the other possibility to help phase lock acquisition, Rana suggested to use a commercial discriminator from Miteq.


(principle idea)

The diagram below shows a schematic of the circuit which Matt has built.

FD.png

Basically an input signal is split into two signals right after the input, then one signal goes through directly to a NAND comparator.

On the other hand another split signal goes through a delay line which composed by some RC filters, then arrive at the NAND comparator with a certain amount of delay.

After going through the NAND comparator, the signal looks like a periodic pulses (see below).

If we put a signal of higher frequency we get more number of pulses after passing through the NAND.

pulses.png

Finally the pulse-signal will be integrated at the low pass filter and converted to a DC signal.

Thus the amplitude of DC signal depends on the number of the pulses per unit time, so that the output DC signal is proportional to the frequency of an input signal.

 

 

(result)

By putting a TTL high-low signal, an output of the circuit shows 100kHz/V linear response.

It means we can get DC voltage of 1 V if a signal of 100kHz is injected into the input.

And the noise measurement has been done while injecting a input signal. The noise level of 0.2Hz/rtHz @ 10 Hz was yielded.

Therefore we can lock the green PLL by using an ordinary VCO loop after we roughly guide a beat note by using this kind of discriminator.

 FDnoise.png

Attachment 1: DSC_1407.JPG
DSC_1407.JPG
Attachment 2: FD.png
FD.png
Attachment 3: FDnoise.png
FDnoise.png
  2717   Sat Mar 27 16:23:10 2010 KojiUpdateSUSanother SRM sidemagnet glued

Kiwamu and Koji

Last night we have released PRM from the gluing fixture. All of the six magnets are successfully released from the fixture.

We put SRM on the fixuture and glued a side magnet which we had failed at the last gluing.

We let it cure in the Al house. This should be the last magnet gluing until ETMs are delivered.

[Current status]

ITMX (ITMU03): all of magnets/guiderod/standoffs glued, mirror baked; balance to be confirmed
ITMY (ITMU04): all of magnets/guiderod/standoffs glued, balance confirmed, mirror baked
SRM  (SRMU03): magnets/guiderod/standoff glued; a side magnet gluing in process, balance to be confirmed, last stand off to be glued, mirror to be baked
PRM  (SRMU04): magnets/guiderod/standoff glued; balance to be confirmed, last stand off to be glued, mirror to be baked

TT:            magnets/guiderod/standoff glued; balance to be confirmed, last stand off to be glued, mirror to be baked

  2716   Fri Mar 26 10:48:35 2010 steveUpdateSAFETYsafety glasses measured

Safety glasses 1064 nm transmission measured at ~200 mW level. They are all good.

Attachment 1: P1060188.JPG
P1060188.JPG
Attachment 2: P1060199.JPG
P1060199.JPG
  2715   Thu Mar 25 17:32:42 2010 AlbertoUpdate40m UpgradingREFL55 Upgraded

I upgraded the old REFL199 to the new REFL55.

To do that I had to replace the old photodiode inside, switching to a 2mm one.

Electronics and optical transfer functions, non normalized are shown in the attached plot.

2010-03-25_REFL55_model_to_meas_comparison.png

The details about the modifications are contained in this dedicated wiki page (Upgrade_09 / RF System / Upgraded RF Photodiodes)

Attachment 1: 2010-03-25_REFL55_model_to_meas_comparison.png
2010-03-25_REFL55_model_to_meas_comparison.png
  2714   Thu Mar 25 17:29:48 2010 kiwamu, mottUpdateGreen LockingPLL two NPROs

In this afternoon, Mott and I tried to find a beat note between two NPROs which are going to be set onto each end table for green locking.

At first time we could not find any beats. However Koji found that the current of innolight NPRO was set to half of the nominal.

Then we increased the current to the nominal of 2A, finally we succeeded in finding a beat note.

Now we are trying to lock the PLL.

P.S. we also succeeded in acquiring the lock

 

nominal setup
  innolight  lightwave
T [deg] 39.75 37.27
current [A] 2 2
laser power [mW] 950 700


  2713   Thu Mar 25 09:07:08 2010 steveUpdatePEMGuralp2 x problem is back

Quote:

I went and double-checked and aligned the styrofoam cooler at ~5:00 UTC. It was fine, but we really need a better huddling box. Where's that granite anyway?

Here's the new Huddle Test output. This time I show the X-axis since there's some coherence now below 0.1 Hz.

You'll also notice that the Wiener filter is now beating the FD subtraction. This happened when I increased the # of taps to 8000. Looks like the noise keeps getting lower as I increase the number of taps, but this is really a kind of cheat if you think about it carefully.

 The same thing happening again.  The intermittent offset upstream of the seismometer that never got fixed.

The granite plate and ball bearings are in. I will place seismometers on it.

Attachment 1: grlp2xproblm.jpg
grlp2xproblm.jpg
  2712   Wed Mar 24 15:59:59 2010 kiwamu, mattUpdateGreen Lockingleave PLL locked

Matt checked it in this morning and he found it's been locked during the night.

 

 

  2711   Wed Mar 24 14:57:21 2010 AlbertoUpdate40m UpgradingREFL11 upgraded

 

 Hartmut suggested a possible explanation for the way the electronics transfer function starts picking up at ~50MHz. He said that the 10KOhm resistance in series with the Test Input connector of the box might have some parasitic capacitance that at high frequency lowers the input impedance.

Although Hartmut also admitted that considering the high frequency at which the effect is observed, anything can be happening with the electronics inside of the box.

  2710   Wed Mar 24 14:52:02 2010 AlbertoUpdateGreen Lockingtwo NPRO PLL

Quote:

Now some pedestals, mirrors and lenses are left on the PSL table, since we are on the middle way to construct a PLL setup which employs two NPROs instead of use of PSL laser.

So Please Don't steal any of them.

 Can I please get the network analyzer back?

  2709   Wed Mar 24 12:40:25 2010 daisukeConfigurationGeneralPeriscope for green laser delivery from the BSC to PSL table

The periscope design for beam elevation of the green beams is posted. The design for the 90 deg steering version is also coming.

(2010-03-29: update drawings by daisuke)

90deg version: http://nodus.ligo.caltech.edu:8080/40m/2725

40m_periscope.png

Attachment 2: 40m_periscope_A_100329.pdf
40m_periscope_A_100329.pdf 40m_periscope_A_100329.pdf 40m_periscope_A_100329.pdf 40m_periscope_A_100329.pdf
Attachment 3: 40m_periscope_A_dwg_100329.zip
  2708   Wed Mar 24 12:38:17 2010 HartmutConfigurationGreen LockingBroadband PD for green PLL

Modified one of the PD assemblies carrying a large SI-Diode (~10mm diameter).

Removed elements used for resonant operation and changed PD readout to transimpedance

configuration. The opamp is a CLC409 with 240 Ohm feedback (i.e. transimpedance) resistor.

To prevent noise peaking at very high frequencies and get some decoupling of the PD,

I added a small series resistor in line with the PD and the inverting opamp input.

It was chosen as 13 Ohm, and still allows for operation up to ~100MHz.

Perhaps it could be smaller, but much more bandwith seems not possible with this opamp anyway.

Changes are marked in the schematic, and I list affected components here.

(Numbers refer to version 'PD327.SCH' from 30-April-1997):

-removed L4

-connected L3 (now open pad) via 100 Ohm to RF opamp output. This restores the DC sognal output.

-removed c17

-connected pin 3 of opamp via 25 Ohm to GND

-connected kathode of PD via 13 Ohm to pin 2 of opamp

-removed L6, C26, L5, C18, and C27

-shorted C27 pad to get signal to the RF output

 

Measured the optical TF with the test laser setup.

(Note that this is at 1064nm, although the PD is meant to work with green light at 532nm!)

Essentially it looks usable out to 100MHz, where the gain dropped only by about

6dB compared to 10MHz.

Beyond 100MHz the TF falls pretty steeply then, probably dominated by the opamp.

 

The maximal bias used is -150V.

If the bias is 'reduced' from -150V to -50V, the response goes down by 4dB at 10MHz and

by 9dB at 100MHz.

 The average output was 30mV at the RF output, corresponding to 60mV at the opamp output (50Ohm divider chain).

With 240 Ohm transimpedance this yields 250µA photo-current used for these transfer functions.

SiAmpl.png

 

SiPhase.png

 

 

  2707   Wed Mar 24 04:22:51 2010 kiwamu, mattUpdateGreen Lockingtwo NPRO PLL

Now some pedestals, mirrors and lenses are left on the PSL table, since we are on the middle way to construct a PLL setup which employs two NPROs instead of use of PSL laser.

So Please Don't steal any of them.

  2706   Wed Mar 24 03:58:18 2010 kiwamu, matt, kojiUpdateGreen Lockingleave PLL locked

We are leaving the PLL as it is locked in order to see the long term stability. And we will check the results in early morning of tomorrow.

DO NOT disturb our PLL !!

  


(what we did)

After Mott left, Matt and I started to put feedback signals to the temperature control of NPRO.

During doing some trials Matt found that NPRO temperature control input has an input resistance of 10kOhm.

Then we put a flat filter ( just a voltage divider made by a resistor of ~300kOhm and the input impedance ) with a gain of 0.03 for the temperature control to inject a relatively small signal, and we could get the lock with the pzt feedback and it.

In addition, to obtain more stable lock we then also tried to put an integration filter which can have more gain below 0.5Hz.

After some iterations we finally made a right filter which is shown in the attached picture and succeeded in obtaining stable lock.

 

 

 

Attachment 1: DSC_1402.JPG
DSC_1402.JPG
  2705   Wed Mar 24 02:06:24 2010 KojiUpdateIOOvac envelope has to be sealed as antproof for overnight

Matt and Koji:

We closed the light doors of the chambers.

Quote:

Roger.

Quote:

 This seal is good for daily use- operation only. The IFO has to be sealed  with light metal doors every night so ants and other bugs can not find their way in.

 

 

  2704   Tue Mar 23 22:46:43 2010 AlbertoUpdate40m UpgradingREFL11 upgraded
I modified REFL11 according to the changes lsited in this schematic (see wiki  / Upgrade 09 / RF System / Upgraded RF Photodiodes ).
I tuned it to be resonant at 11.06MHz and to have a notch at 22.12MHz.
These are the transfer functions that I measured compared with what I expected from the LISO model.

2010-03-23_REFL11_model_to_meas_comparison.png

The electronics transfer function is measured directily between the "Test Input" and the "RF Out" connector of the box. the optical transfer function is measured by means of a AM laser (the "Jenne laser") modulated by the network analyzer.
The AM laser's current was set at 20.0mA and the DC output of the photodiode box read about 40mV.
The LISO model has a different overall gain compared to the measured one, probably because it does not include the rest of the parts of the circuit other than the RF out path.

I spent some time trying to understand how touching the metal cage inside or bending the PCB board affected the photodiode response. It turned out that there was some weak soldering of one of the inductors.

  2703   Tue Mar 23 18:44:46 2010 MottUpdateABSLPLL reconstructed

 

 After realigning and getting the lock today, I tried to add in the SR785 to measure the transfer function.  As soon as I turn on the piezo input on the PDH box, however, the lock breaks and I cannot reacquire it.  We are using an SR650 to add in the signal from the network analyzer and that has worked. We also swapped the 20 dB attenuator for a box which mimics the boost functionality (-20 dB above 100 Hz, 0 dB below 6Hz).  I took some spectra with the SR750, and will get some more with the network analyzer once Alberto has finished with it. 

The SR750 spectra is posted below.  The SR750 only goes up to 100 kHz, so I will have to use the network analyzer to get the full range. 

Attachment 1: NPRO_PLL_freqresp.png
NPRO_PLL_freqresp.png
  2702   Tue Mar 23 15:38:26 2010 AlbertoUpdateelogelog just restarted

I found the elog down and I restarted it.

Then, after few seconds it was down again. Maybe someone else was messing with it. I restarted an other 5 times and eventually it came back up.

  2701   Tue Mar 23 10:11:13 2010 steveUpdatePEM small air cond unit failed yesterday

Old control room air condition failed yesterday around noon. It was blowing 80-85F hot air for about 2-3 hours at racks 1Y4-7 and the entry room 103

Attachment 1: 20datm.jpg
20datm.jpg
  2700   Tue Mar 23 09:55:20 2010 KojiUpdateIOOvac envelope has to be sealed as antproof for overnight

Roger.

Quote:

 This seal is good for daily use- operation only. The IFO has to be sealed  with light metal doors every night so ants and other bugs can not find their way in.

 

  2699   Tue Mar 23 09:37:36 2010 steveUpdateIOOvac envelope has to be sealed as antproof for overnight

Quote:

This is the first touch to the MC mirrors after the earthquake on 16th.

  • I made an aluminum access connector so that we can work on the MC even the door is open. We still can be able to open the aluminum tube. The photos are attached. Steve, could you please look it at a glance whether the seal is enough or not.
  • MC resonances were flashing. Align MC2 and MC3 so that we have many TEM00s.
  • Found c1vmesus2 gone mad. Restarted remotely according to the wiki entry. 
  • Reset the MC coil output matrix to 1. (Previously, balance was adjusted so that A2L was minimized.)
  • Excite MC2 Pitch/Yaw at 8 and 9 Hz, looking at the peaks in the MC-MCL output. Move MC2 Pitch/Yaw so that the peak
    is reduced. (*)
  • MC1/MC3 were aligned so that we get the maximum transmission (or minimum reflection). (**)
  • Repeat (*) and (**)

So far, I have aligned in Yaw such that the yaw peak is minimized.

 This seal is good for daily use- operation only. The IFO has to be sealed  with light metal doors every night so ants and other bugs can not find their way in.

Our janitor Kevin is mopping the hole IFO room floor area with 5%  ant killing solution in water in order to discourage bugs getting close to our openings of the vented chamber.

You may be sensitive to this chemical too.  Do not open chamber till after lunch.

Attachment 1: pc3.JPG
pc3.JPG
Attachment 2: pc4.JPG
pc4.JPG
  2698   Tue Mar 23 00:31:51 2010 KojiUpdateIOOMC realigned

This is the first touch to the MC mirrors after the earthquake on 16th.

  • I made an aluminum access connector so that we can work on the MC even the door is open. We still can be able to open the aluminum tube. The photos are attached. Steve, could you please look it at a glance whether the seal is enough or not.
  • MC resonances were flashing. Align MC2 and MC3 so that we have many TEM00s.
  • Found c1vmesus2 gone mad. Restarted remotely according to the wiki entry. 
  • Reset the MC coil output matrix to 1. (Previously, balance was adjusted so that A2L was minimized.)
  • Excite MC2 Pitch/Yaw at 8 and 9 Hz, looking at the peaks in the MC-MCL output. Move MC2 Pitch/Yaw so that the peak
    is reduced. (*)
  • MC1/MC3 were aligned so that we get the maximum transmission (or minimum reflection). (**)
  • Repeat (*) and (**)

So far, I have aligned in Yaw such that the yaw peak is minimized.

Attachment 1: IMG_2346.jpg
IMG_2346.jpg
Attachment 2: IMG_2347.jpg
IMG_2347.jpg
  2697   Mon Mar 22 23:37:32 2010 MottUpdateABSLPLL reconstructed

 

We have managed to lock the PLL to reasonable stability. The RF input is attenuated by 26 dBm and the beat signal locks very close to the carrier of the marconi (the steps on the markers of the spectrum analyzer are coarse).  We can use the marconi and the local boost of the pdh box to catch the lock at 0 gain.  Once the lock is on, the gain can be increased to stabilize the lock.  The locked signals are shown in the first photo (the yellow is the output of the mixer and the blue is the output to the fast input of the laser.  If the gain is increased too high, the error signal enters an oscillatory regime, which probably indicates we are overloading the piezo.  This is shown in the second photo, the gain is being increased in time and we enter the non-constant regime around mid-way through.

Tomorrow I will use this locked system to measure the PZT response (finally!).

Attachment 1: 2010-03-22_23.14.00.jpg
2010-03-22_23.14.00.jpg
Attachment 2: 2010-03-22_23.24.50.jpg
2010-03-22_23.24.50.jpg
  2696   Mon Mar 22 22:11:26 2010 MottUpdateABSLPLL reconstructed

 

 It looks like the PLL drifted alot over the weekend, and we couldn't get it back to 9 dBm.  We switched back to the new focus wideband PD to make it easier to find the beat signal.  I replaced all the electronics with the newly fixed UPDH box (#17) and we aligned it to the biggest beat frequency we could get, which ended up being -27 dBm with a -6.3V DC signal from the PD.  

Locking was still elusive, so we calculated the loop gain and found the UGF is about 45 kHz, which is too high.  We added a 20 dB attenuator to the RF input to suppress the gain and we think we may have locked at 0 gain.  I am going to add another attenuator (~6 dB) so that we can tune the gain using the gain knob on the UPDH box.  

Finally, attached is a picture of the cable that served as the smb - BNC adaptor for the DC output of the PD.  The signal was dependent on the angle of the cable into the scope or multimeter.  It has been destroyed so that it can never harm another innocent experiment again!

Attachment 1: IMG_0150.JPG
IMG_0150.JPG
ELOG V3.1.3-